Using Laplace Transform (LT) and Inverse Laplace Transform (LT) solve the following system of equations: 1. X=- 2x + 5oy y
Answer #1

Doubt in this then comment below.. i will help you..

By rules and regulations we are allow to do only one question at a time..so i do only 1st

.

please thumbs up for this solution...thanks..

.

x + 450 25 25 25 Laploco Apply syrs) - 251 als) - 2x1s) y+ y 25 (st 4r5)= o 25 (s+2) 50 25/543) 258+10 25 25225 2(s+ 25 Laplace invese Apply -t/25 St/25 175 pow (St 25 4rs1 = 4(57_ Laplo o inveese Aply -t/25 -3+/25 yltls 25